Hi there i have two questions, i'm in geometry and i have two questions on my homework could you help?
the directions are:
find the length and with of the rectangle described
the two of the rectangles are described like
1.) the length is 12 more than the width. the perimeter is 7 times the width
2.) the length is 4 units less than 3 times the width. The perimeter is 22 units more than twice the width
these are two of the 15 problems i have to do for homework any help will be apprieciated
Thank You

Answer by edjones(8007)

1)
L=w+12
P=7w
2L+2w=7w
2(w+12)+2w=7w Substitute w+12 for L
2w+24+2w-7w=0
-3w=-24
w=8
L=20
.
Ed